### v1

@@ Appendix A Probabilistic construction of multiple rank-1 lattices
For $\alpha>1/2$, $d\in\mathbb{N}$, a collection of non-negative weights $\boldsymbol{\gamma}=(\gamma_{u})_{u\subset\mathbb{N},|u|<\infty}$, and $M>0$, let $\mathcal{A}_{d}$ be defined as in (2). Under the same assumptions as in Proposition A.1, for any $\lambda>1/\alpha$ and $\beta\in(0,1)$, there exist a constant $C_{c,\delta,\beta}>0$ such that
@@ Appendix B Lower bounds for single lattice-based algorithms
In this appendix, we provide two approaches to derive lower bounds on the worst-case $L_{2}$ error for single lattice-based algorithms. For simplicity, we consider the two-dimensional unweighted Korobov space with smoothness $\alpha>1/2$, and the number of points $N$ is assumed to be a prime number. The essence of both approaches lies in identifying a short non-zero vector in the dual lattice (which aliases to the origin) to construct a fooling function such that the corresponding approximation error is as large as possible. The first approach, relying on the pigeonhole principle, is based on [10, Theorem 9.15]. This result is attributed to Korobov [16] and was recently exploited in [3] (without referring to earlier works). The second approach, which is constructive and relies on the properties of continued fractions, follows the argument in [10, Theorem 9.14], though we present a simplified derivation for clarity.
B.1 A lower bound by Korobov
@@ B.2 A lower bound by Hua and Wang
For a real number $x$, we denote its continued fraction convergents by $p_{t}/q_{t}$. It is well-known that

### v2

@@ Appendix A Probabilistic construction of multiple rank-1 lattices
For $\alpha>1/2$, $d\in\mathbb{N}$, a collection of non-negative weights $\boldsymbol{\gamma}=(\gamma_{u})_{u\subset\mathbb{N},|u|<\infty}$, and $M>0$, let $\mathcal{A}_{d}$ be defined as in (2). Under the same assumptions as in Proposition A.1, for any $\lambda>1/\alpha$ and $\beta\in(0,1)$, there exists a constant $C_{c,\delta,\beta}>0$ such that
@@ Appendix B Lower bounds for single lattice-based algorithms
In this appendix, we provide two approaches to derive lower bounds on the worst-case $L_{2}$ error for single lattice-based algorithms. For simplicity, we consider the two-dimensional unweighted Korobov space with smoothness $\alpha>1/2$, and the number of points $N$ is assumed to be a prime number. The essence of both approaches lies in identifying a short non-zero vector in the dual lattice (which aliases to the origin) to construct a fooling function such that the corresponding approximation error is as large as possible. The first approach, relying on the pigeonhole principle, is based on [10, Theorem 9.15]. This result can be traced back to the early works of Smolyak [27] and Korobov [16], and was recently exploited in [3] (without referring to earlier works). The second approach, which is constructive and relies on the properties of continued fractions, follows the argument in [10, Theorem 9.14], though we present a simplified derivation for clarity.
B.1 A lower bound by Smolyak and Korobov
@@ B.2 A lower bound by Hua and Wang
To briefly introduce the concept, any real number $x$ can be expanded into a continued fraction of the form $x=[a_{0};a_{1},a_{2},\ldots]$, where $a_{0}$ is an integer and $a_{1},a_{2},\ldots$ are positive integers. Truncating this expansion at $a_{t}$ yields a rational approximation $p_{t}/q_{t}\coloneqq[a_{0};a_{1},\ldots,a_{t}]$, which is called the $t$-th convergent of $x$. It is well-known that